here is what the question is.
Kristen bought her lunch for $6.85. She paid for it with nickels and quarters. If there were 53 quarters, how many of each coin type was there?

Answer by Alan3354(69443)

53 quarters = $13.25 ???
Do you mean she used 53 of them? Or they were just in the vicinity?
===========
53 nickels = 265 cents.
Each nickel replaced by a quarter add 20 cents
----
(685 - 265)/20 = 420/21 = 21 quarters
--> 21 quarters & 32 nickels